6x+2y=18 find the slope and y intercept
Answer by Simnepi(216)

The form y = mx + c gives the equation of a straight line, where m is the slope of the line(the gradient) and b is where it crosses the y axis (this is known as the intercept of the line and the y axis, or intercept).
All lines with the same gradient (i.e. the same number in front of the x) are parallel.
First, let's rearrange your line equation into the correct form of y = mx + c
6x+2y=18
We move the 6x across the equals sign, changing its sign as we do so
So now we have
2y=-6x+18
Now we divide by 2 to get rid of the 2 in front of the y term (remembering to divide EVERYTHING by 2, not just the 2y!) giving
y=-3x+9
Now we can say the slope is -3 and the intercept is 9.